# The spectral continuum in the Rabi-Stark model

### v1

@@ 2 Effective quantum Rabi model
In the Bargmann representation there are two requirements for a function $\psi(z)$ to be an element of ${\cal B}$. The first Bargmann condition requires analyticity of $\psi(z)$ in all of $\mathbb{C}$, $\psi(z)$ must be entire. The second Bargmann condition reads $\langle\psi|\psi\rangle<\infty$. Both conditions will play a crucial role in the exact determination of the spectrum of $H_{RS}$ given in Equ. (3).
@@
gets represented by the coupled system of first-order ODE’s [35]
@@
For $\gamma=0$, (14,15) is equivalent to (9,10). Division by $\omega+\gamma$ leads to the familiar form [4]
@@
We note that the coupling constant $g$ is renormalized by the factor $(\omega^{2}-\gamma^{2})^{-1/2}$ which diverges for $\gamma\rightarrow\omega_{-}$, although less so than the renormalization of energy $E$ and qubit splitting $\Delta$, which are mixed for $\gamma\neq 0$. Using the manifest $Z_{2}$ invariance of (18,19) and adopting the G-function formalism developed by one of us in [4], we obtain the $G$-functions
@@
The distance between the poles at $E^{(p)}_{n}$ and $E^{(p)}_{n-1}$ corresponding to integer $x$ is $\delta E=(\omega^{2}-\gamma^{2})/\omega$, in accord with the results in [36, 37, 38]. This distance tends to zero for $\gamma\rightarrow\omega_{-}$, when the Stark-coupling $\gamma$ becomes critical, similar to the two-photon QRM [17] for $g=\omega/2$. In the latter case the spectrum develops a continuous part at the critical coupling. This continuous part of the spectrum could be inferred from the $G$-conjecture [4], which assumes that the distribution of zeroes of $G_{\pm}(E)$ is to some extend determined by its poles. It reads,
$\bm{G}$-conjecture: Let $\{W_{j}|j\in\rm{I\!N}\}$ be the set of (simple) poles of $G(E)$. The zeroes of $G(E)$, forming the regular spectrum, $\{E_{k}|k\in\rm{I\!N}\}$ are either located below the first pole, $E_{k}<W_{0}$, or between two poles, $W_{j}<E_{k}<W_{j+1}$. In the latter case, there are at most two such zeros in $]W_{j},W_{j+1}[$, i.e. $W_{j}<E_{k_{l}}<E_{k_{l}+1}<W_{j+1}$. If this is the case, the intervals $]W_{j-1},W_{j}[$ and $]W_{j+1},W_{j+2}[$ may contain either none or only one $E_{k}$. If the interval $]W_{j},W_{j+1}[$ contains none of the $E_{k}$, the adjacent intervals $]W_{j-1},W_{j}[$ and $]W_{j+1},W_{j+2}[$ contain at least one $E_{k}$.
This conjecture has been seen to be valid by extensive numerical checks in case of the QRM [57]. Very recently, the conjecture has been proven to be valid for almost all intervals $[]W_{j},W_{j+1}[$58]. The conjecture entails that at least one zero of $G_{\pm}(E)$ is located between $W_{j}$ and $W_{j+2}$. If the distance between the poles approaches zero as one of the parameters approaches a critical value, the discrete eigenvalues become more and more dense on the real axis above $W_{0}$, eventually forming a continuum if the critical point is reached.
@@
From (21) we see that $G_{+}(x)=G_{-}(x)$ for $\tilde{\Delta}=0$ or $E=-\omega\Delta/\gamma$. If this energy happens to be an eigenvalue, it must be doubly degenerate. Indeed, the case $\tilde{\Delta}=0$ corresponds to a degenerate qubit in the renormalized model, therefore $\tilde{E}+\tilde{g}^{2}=n$ with $n=0,1,2,\ldots$ which leads to the array of special crossing points at the same energy $E=-\omega\Delta/\gamma$ independent from $g$ in the spectral graph. Compare Fig.2 in [38] and Equation (4) above [^4] The notation in [38] corresponds to the replacement $\gamma\rightarrow U/2$ and $\Delta\rightarrow\Delta/2$.. In contrast to the ordinary QRM, there is a crossing on the first baseline $x=n=0$, corresponding to $E=-(\gamma\Delta+g^{2})/\omega$, which translates to the condition on the parameters $g^{2}=\Delta(\omega^{2}-\gamma^{2})/\gamma$, which can be satisfied for $\gamma>0$ (if $\Delta>0$ as usual) [38]. If $\Delta$ is sufficiently small, this crossing corresponds to a degeneracy of the ground state with the first excited state and thus to a quantum phase transition of first order, as pointed out in [38]. We observe that the location of this phase transition will occur for small coupling $g$ (within the reach of experimental platforms) if $\gamma$ is close enough to $\omega$.
3 The limit $\bm{\gamma\rightarrow\omega_{-}}$
We consider the limit $\gamma\rightarrow\omega_{-}$ and start with Equations (18,19). For $\gamma<\omega$, the two regular singular points of this system are located at $z=\pm\tilde{g}=g/\sqrt{\omega^{2}-\gamma^{2}}$. If $\gamma$ approaches $\omega$ from below, these singular points tend to infinity to coalesce at $z=\infty$ for $\gamma=\omega$, forming there an irregular singular point of s-rank three [59]. In terms of the ODE’s (18,19), this process is called confluence of the two singularities and well known e.g. from the hypergeometric equation, where the confluence of two regular singular points at infinity produces the Bessel equation [59]. The first step in deriving the confluent equation is elimination of $\bar{\phi}_{2}$ in terms of $\bar{\phi}_{1}$. We apply the operator on the left hand side of (19) to (18) and obtain
@@ 3 The limit $\bm{\gamma\rightarrow\omega_{-}}$
The parameters in (23) diverge as a power of $\epsilon^{-1}$ with $\epsilon=\sqrt{\omega^{2}-\gamma^{2}}$. For example, $\tilde{g}^{2}$ diverges as $\epsilon^{-2}$, as do $\tilde{E}$ and $\tilde{\Delta}$. The maximally divergent terms in (23) seem to be $\tilde{E}^{2}$ and $\tilde{\Delta}^{2}$, diverging like $\epsilon^{-4}$. The confluence process consists now in letting $\epsilon\rightarrow 0$ while keeping the most divergent terms. In our case these are actually $\sim\epsilon^{-2}$ because
@@
The form of (26) shows it is the eigenvalue problem of an operator which can be written as linear combination of the elements of $\mathfrak{sl}_{2}(\mathbb{R})$, acting in ${\cal B}$ as [60],
@@
and it depends on $\beta_{2}$ and $\rho$, whether $\bar{\phi}_{1}(z)$ is normalizable or not [17]. In our case, we find for $\beta_{2}$ the equation,
@@
which entails $\beta_{2}=0$ for $\alpha\neq\pm 1$. Exactly at $\alpha=\pm 1$, the exponents $\beta_{1}$ become degenerate, either $\beta^{\pm}_{1}=-1$ or $\beta^{\pm}_{1}=1$, which hints at the special role of this case. $E=-\Delta$ ($\beta_{1}^{\pm}=-1$) is the limit of $E=-\omega\Delta/\gamma$, i.e. energy corrsponding to the array of special solutions mentioned in section 2. For $|\alpha|<1$, the normalizability of $\bar{\phi}_{1}(z)$ is determined by $\rho$,
@@
independent from the value of $\beta_{1}$. The fundamental property of asymptotic expansions like (30) is the apperance of the Stokes phenomenon: A certain combination of exponents $\{\beta_{1},\beta_{2},\rho\}$ is only valid in a certain sector of the complex plane, centered around the (Stokes) ray $z(t)=te^{i\psi}$ where $\bar{\phi}_{1}(z)$ is either dominant or recessive [59]. In general, the analytic continuation of a solution having an asymptotic expansion with exponents $\{\beta_{1},\beta_{2},\rho\}$ in a given sector will have different exponents in the other sectors. In our case, the Stokes rays have the form $z=\pm|z|e^{-i\theta/2}$ for $\beta_{1}=e^{i\theta}$ if the solution has dominant asymptotics, $\bar{\phi}_{1}(z)\sim\exp(|z|^{2}/2)$, in the sector $-\theta/2-\pi/4<\textrm{arg}(z)<-\theta/2+\pi/4$. Dominant asymptotics means in the present case, $|\beta_{1}|=1$, that the Bargmann norm of $\bar{\phi}_{1}$ is finite if and only if the integral
@@
is finite [17]. Here we have used the unitary transformation $z\rightarrow e^{i\theta/2}z$ to set $\beta_{1}=1$, which means that the critical sectors, where this case may occur, are $S_{0}=-\pi/4<\textrm{arg}(z)<\pi/4$ and $S_{3}=3\pi/4<\textrm{arg}(z)<5\pi/4$ ($I_{0}(R)$ refers to $S_{0}$). The parameter $R$ must be large enough to ensure validity of the asymptotic expansion in the given sector to the supposed precision. We have ${\rm Re}(\rho)=-1/2$ which is, according to the analysis in [17], the marginal case. Marginal means here that the solution $\bar{\phi}_{1}(z)$ is normalizable if ${\rm Re}(\rho)<-1/2$ and not normalizable for ${\rm Re}(\rho)>-1/2$ [17]. To see this, let’s look at the integration over $y$ in (36),
@@
This integral converges for ${\rm Re}(\rho)<-1/2$ and diverges logarithmically for ${\rm Re}(\rho)=-1/2$. The formal solution $\bar{\phi}_{1}(z)$ has no finite Bargmann norm for $|\alpha|<1$, in accord with the result in [35]. There are no isolated eigenvalues in the energy range $-\Delta-2g^{2}/\omega<E<-\Delta$. Now we have two options:
@@
To decide between these two possibilities, it is not enough to state merely the non-normalizability of the entire formal solutions of (26) as in [35] because these formal solutions are entire (and non-normalizable) also in the gaps between the discrete eigenvalues for $|\alpha|>1$.
To show that there is indeed a continuum, we must demonstrate that the formal eigenfunctions $\bar{\phi}_{\Lambda}(z)$ for eigenvalue $\Lambda$ furnish a spectral measure on the real line absolutely continuous with respect to Lebesgue measure [61]. For example, the formal eigenfunctions $\phi_{p}(x)=(2\pi)^{-1/2}e^{ipx}$ of the momentum operator $-i{\rm{d}}/{\rm{d}}x$ satisfy the generalized orthonormality relation,
@@
and furnish thus a proper spectral measure on the real line $-\infty<p<\infty$, because any element of $L^{2}(\mathbb{R})$ can be expanded in terms of the set $\{\phi_{p}(x)\}$ as a Fourier integral. In ${\cal B}$, the functions $\phi_{p}$ read [17],
@@
$\langle\bar{\phi}_{\Lambda_{1}}|\bar{\phi}_{\Lambda_{2}}\rangle\propto$
$\int_{S_{0}}{\rm{d}}x{\rm{d}}y\ e^{-x^{2}-y^{2}}e^{\frac{1}{2}\Big{(}(x+iy)^{2}+(x-iy)^{2}\Big{)}}(x-iy)^{\rho^{\ast}(\Lambda_{1})}(x+iy)^{\rho(\Lambda_{2})}$ (42)
@@
$=C\int_{-\infty}^{\infty}{\rm{d}}t\ \exp\left(it\left[\frac{\Lambda_{1}}{\sqrt{1-\alpha_{1}^{2}}}-\frac{\Lambda_{2}}{\sqrt{1-\alpha_{2}^{2}}}\right]\right)$
@@
with $a=\textrm{ch}\theta-\textrm{sh}\theta$ and $\textrm{ch}\theta=\alpha$. It follows with $a<1$ that $U$ is an element of $SU(1,1)$, a real subgroup of $SL_{2}(\mathbb{C})$ which acts on ${\cal B}$ as isometries [17]. This means that there is a unitary transformation in ${\cal B}$ which maps the left hand side of (29) onto $\textrm{sh}\theta K_{0}$ which is the Hamiltonian of the harmonic oscillator with frequency $\textrm{sh}\theta$ (including the vacuum term). It follows at once that the eigenvalue equation (29) is equivalent to
@@
in accord with [35, 38]. The energy $E$ is given implicitly in (46) and leads to a lower bound for the discrete spectrum with $\alpha>1$, namely $E>$max$(-\Delta,\Delta-\omega)$. In the case $\alpha<-1$, the same reasoning must be slightly modified because now $\textrm{ch}\theta=-\alpha$ and the transformation $U$ which yields the diagonal form in (45) has now $a=\textrm{sh}\theta+\textrm{ch}\theta>1$ and is not an element of $SU(1,1)$. If one exchanges the two eigenvectors, one obtains
@@
again in accord with [35, 38]. Beyond the coupling regime which supports a lower discrete spectrum, the continuum threshold at $E_{thr}=-(\Delta+2g^{2}/\omega)$ is at the same time the ground state energy. As $E_{thr}<E_{c}$, the lower limit of the continuum deduced from the pole structure of the $G$-function, we see that the confluence process going from (23) to (26) is consistent with the continuum predicted from $G_{\pm}(x)$, at least regarding the energy region $E<-\Delta$. On the other hand, the $G$-function also predicts that the continuum should extend over the whole real axis from $E_{c}$ upwards. How can this be reconciled with the confluence process described in (23) and (26)?
@@
This representation preserves the canonical commutation relations. We set now $\zeta=\sqrt{2C}/\epsilon$, this means that $\zeta$ diverges as $\epsilon\rightarrow 0$. In this limit, the transformation is merely “pseudo-canonical”, a phenomenon known from systems with infinitely many degrees of freedom [62]. The constant $C$ has the dimension of energy. Performing the limit yields
@@
For any finite energy $E>E_{thr}$, the generalized eigenstate will be centered around $q_{\pm}(E)=\pm\zeta\sqrt{(E-E_{thr})/C}$ which diverges like $1/\epsilon=(\omega-\gamma)^{-1/2}$. The same is true of the two associated spectral measures on any interval $[E_{1},E_{2}]$ with $E_{1}>E_{thr}$ which corresponds to elements of $L^{2}(\mathbb{R})$ supported respectively on the intervals ±$\pm[q_{+}(E_{1}),q_{+}(E_{2})]$. As the normalizable states of (46) have wave functions exponentially decreasing for $q\rightarrow\pm\infty$, we conclude that they will no longer hybridize with the states in (59) for $\gamma\rightarrow\omega_{-}$ and belong to a dynamically invariant subspace of $L^{2}(\mathbb{R})$. Therefore, they can be considered as “bound states in the continuum” at the critical point. However, for all $\gamma<\omega$ the precursors of the BICs will hybridize with the precursors of the states in (59), leading to avoided crossings in the spectral graph which become more and more narrow as $\gamma$ approaches $\omega$. True BICs exists only at the critical point itself, which is here defined as the limit $\gamma\rightarrow\omega_{-}$, not by setting $\gamma=\omega$ in (9) and (10). The latter case was studied in [35, 38] and leads just to the two discrete spectra (46), (48) and the “small continuum”, obtained by the first confluence process given in (26). The canonical transformation (54) can be implemented unitarily only if $\zeta<\infty$, so it fails exactly at $\gamma=\omega$ which is a singular point of the transformation. This phenomenon bears close resemblance to the failure to implement the scale transformation (54) unitarily with $\zeta<\infty$ but for infinitely many bosonic variables [62]. The Rabi-Stark model at the critical point provides thus an analogue of the infrared divergences occurring in interacting quantum field theories – just like the singular limit $\Delta/\omega\rightarrow\infty$ leads to a quantum phase transition of the QRM [11], although the model contains only two degrees of freedom. The crucial difference being that in the latter case the model parameters have to take unphysical (infinite) values, whereas the Rabi-Stark model can be simulated within cavity QED right at the critical point $\gamma=\omega$ [30]. Our results entail that this critical point has to be regularized by setting $\gamma=\omega-\epsilon^{2}$, before letting $\epsilon\rightarrow 0$ to obtain the full spectrum at $\gamma=\omega$.
4 Numerical results for large-dimensional state space

### v2

@@ II Effective quantum Rabi model
In the Bargmann representation there are two requirements for a function $\psi(z)$ to be an element of ${\cal B}$. The first Bargmann condition requires analyticity of $\psi(z)$ in all of $\mathbb{C}$, $\psi(z)$ must be entire. The second Bargmann condition reads $\langle\psi|\psi\rangle<\infty$. Both conditions will play a crucial role in the exact determination of the spectrum of $H_{RS}$ given in Equation (3).
@@
gets represented by the coupled system of first-order ODE’s maciejewski2015
@@
For $\gamma=0$, Equations (14,15) are equivalent to (9,10). Division by $\omega+\gamma$ leads to the familiar form braak2011
@@
We note that the coupling constant $g$ is renormalized by the factor $(\omega^{2}-\gamma^{2})^{-1/2}$ which diverges for $\gamma\rightarrow\omega_{-}$, although less so than the renormalization of energy $E$ and qubit splitting $\Delta$, which are mixed for $\gamma\neq 0$. Using the manifest $\mathbb{Z}_{2}$ invariance of (18,19) and adopting the G-function formalism developed by one of us in braak2011 , we obtain the $G$-functions
@@
The distance between the poles at $E^{(p)}_{n}$ and $E^{(p)}_{n-1}$ corresponding to integer $x$ is $\delta E=(\omega^{2}-\gamma^{2})/\omega$, in accord with the results in eckle2017 ; eckle2023 ; xie2019 . This distance tends to zero for $\gamma\rightarrow\omega_{-}$, when the Stark-coupling $\gamma$ becomes critical, similar to the two-photon QRM braak2023 for $g=\omega/2$. In the latter case the spectrum develops a continuous part at the critical coupling. This continuous part of the spectrum could be inferred from the $G$-conjecture braak2011 , which assumes that the distribution of zeroes of $G_{\pm}(E)$ is to some extend determined by its poles. It reads,
$\bm{G}$-conjecture: Let $\{W_{j}|j\in\rm{I\!N}\}$ be the set of (simple) poles of $G(E)$. The zeroes of $G(E)$, forming the regular spectrum, $\{E_{k}|k\in\rm{I\!N}\}$ are either located below the first pole, $E_{k}<W_{0}$, or between two poles, $W_{j}<E_{k}<W_{j+1}$. In the latter case, there are at most two such zeros in the open interval $]W_{j},W_{j+1}[$, i.e. $W_{j}<E_{k_{l}}<E_{k_{l}+1}<W_{j+1}$. If this is the case, the intervals $]W_{j-1},W_{j}[$ and $]W_{j+1},W_{j+2}[$ may contain either none or only one $E_{k}$. If the interval $]W_{j},W_{j+1}[$ contains none of the $E_{k}$, the adjacent intervals $]W_{j-1},W_{j}[$ and $]W_{j+1},W_{j+2}[$ contain at least one $E_{k}$.
This conjecture has been seen to be valid by extensive numerical checks in case of the QRM linh2024 . Very recently, the conjecture has been proven to be valid for almost all intervals $]W_{j},W_{j+1}[$ rudnick2023 . The conjecture entails that at least one zero of $G_{\pm}(E)$ is located between $W_{j}$ and $W_{j+2}$. If the distance between the poles approaches zero as one of the parameters approaches a critical value, the discrete eigenvalues become more and more dense on the real axis above $W_{0}$, eventually forming a continuum if the critical point is reached.
@@
From Equation (21) we see that $G_{+}(x)=G_{-}(x)$ for $\tilde{\Delta}=0$ or $E=-\omega\Delta/\gamma$. If this energy happens to be an eigenvalue, it must be doubly degenerate. Indeed, the case $\tilde{\Delta}=0$ corresponds to a degenerate qubit in the renormalized model, therefore $\tilde{E}+\tilde{g}^{2}=n$ with $n=0,1,2,\ldots$ which leads to the array of special crossing points at the same energy $E=-\omega\Delta/\gamma$ independent from $g$ in the spectral graph. Compare Fig.2 in xie2019 and Equation (4) above note3 . In contrast to the ordinary QRM, there is a crossing on the first baseline $x=n=0$, corresponding to $E=-(\gamma\Delta+g^{2})/\omega$, which translates to the condition on the parameters $g^{2}=\Delta(\omega^{2}-\gamma^{2})/\gamma$, which can be satisfied for $\gamma>0$ (if $\Delta>0$ as usual) xie2019 . If $\Delta$ is sufficiently small, this crossing corresponds to a degeneracy of the ground state with the first excited state and thus to a quantum phase transition of first order, as pointed out in xie2019 . We observe that the location of this phase transition will occur for small coupling $g$ (within the reach of experimental platforms) if $\gamma$ is close enough to $\omega$.
III The limit $\bm{\gamma\rightarrow\omega_{-}}$
We consider the limit $\gamma\rightarrow\omega_{-}$ and start with Equations (18,19). For $\gamma<\omega$, the two regular singular points of this system are located at $z=\pm\tilde{g}=g/\sqrt{\omega^{2}-\gamma^{2}}$. If $\gamma$ approaches $\omega$ from below, these singular points tend to infinity to coalesce at $z=\infty$ for $\gamma=\omega$, forming there an irregular singular point of s-rank three slavyanov2000 . In terms of the ODE’s (18,19), this process is called confluence of the two singularities and well known e.g. from the hypergeometric equation, where the confluence of two regular singular points at infinity produces the Bessel equation slavyanov2000 . The first step in deriving the confluent equation is elimination of $\bar{\phi}_{2}$ in terms of $\bar{\phi}_{1}$. We apply the operator on the left hand side of (19) to (18) and obtain
@@ III The limit $\bm{\gamma\rightarrow\omega_{-}}$
The parameters in (23) diverge as a power of $\epsilon^{-1}$ with $\epsilon=\sqrt{\omega^{2}-\gamma^{2}}$. For example, $\tilde{g}^{2}$ diverges as $\epsilon^{-2}$, as do $\tilde{E}$ and $\tilde{\Delta}$. The maximally divergent terms in Equation (23) seem to be $\tilde{E}^{2}$ and $\tilde{\Delta}^{2}$, diverging like $\epsilon^{-4}$. The confluence process consists now in letting $\epsilon\rightarrow 0$ while keeping the most divergent terms. In our case these are actually $\sim\epsilon^{-2}$ because
@@
The form of Equation (26) shows it is the eigenvalue problem of an operator which can be written as linear combination of the elements of $\mathfrak{sl}_{2}(\mathbb{R})$, acting in ${\cal B}$ as vourdas2006 ,
@@
and it depends on $\beta_{2}$ and $\rho$, whether $\bar{\phi}_{1}(z)$ is normalizable or not braak2023 . In our case, we find for $\beta_{2}$ the equation,
@@
which entails $\beta_{2}=0$ for $\alpha\neq\pm 1$. Exactly at $\alpha=\pm 1$, the exponents $\beta_{1}$ become degenerate, either $\beta^{\pm}_{1}=-1$ or $\beta^{\pm}_{1}=1$, which hints at the special role of this case. $E=-\Delta$ ($\beta_{1}^{\pm}=-1$) is the limit of $E=-\omega\Delta/\gamma$, the energy corresponding to the array of special solutions mentioned in section II. For $|\alpha|<1$, the normalizability of $\bar{\phi}_{1}(z)$ is determined by $\rho$,
@@
independent from the value of $\beta_{1}$. The fundamental property of asymptotic expansions like (30) is the appearance of the Stokes phenomenon: A certain combination of exponents $\{\beta_{1},\beta_{2},\rho\}$ is only valid in a certain sector of the complex plane, centered around the (Stokes) ray $z(t)=te^{i\psi}$ where $\bar{\phi}_{1}(z)$ is either dominant or recessive slavyanov2000 . In general, the analytic continuation of a solution having an asymptotic expansion with exponents $\{\beta_{1},\beta_{2},\rho\}$ in a given sector will have different exponents in the other sectors. In our case, the Stokes rays have the form $z=\pm|z|e^{-i\theta/2}$ for $\beta_{1}=e^{i\theta}$ if the solution has dominant asymptotics, $\bar{\phi}_{1}(z)\sim\exp(|z|^{2}/2)$, in the sector $-\theta/2-\pi/4<\textrm{arg}(z)<-\theta/2+\pi/4$. Dominant asymptotics means in the present case, $|\beta_{1}|=1$, that the Bargmann norm of $\bar{\phi}_{1}$ is finite if and only if the integral
@@
is finite braak2023 . Here we have used the unitary transformation $z\rightarrow e^{i\theta/2}z$ to set $\beta_{1}=1$, which means that the critical sectors, where this case may occur, are $S_{0}=-\pi/4<\textrm{arg}(z)<\pi/4$ and $S_{3}=3\pi/4<\textrm{arg}(z)<5\pi/4$ ($I_{0}(R)$ refers to $S_{0}$). The parameter $R$ must be large enough to ensure validity of the asymptotic expansion in the given sector to the supposed precision. We have ${\rm Re}(\rho)=-1/2$ which is, according to the analysis in braak2023 , the marginal case. Marginal means here that the solution $\bar{\phi}_{1}(z)$ is normalizable if ${\rm Re}(\rho)<-1/2$ and not normalizable for ${\rm Re}(\rho)>-1/2$ braak2023 . To see this, let’s look at the integration over $y$ in (36),
@@
This integral converges for ${\rm Re}(\rho)<-1/2$ and diverges logarithmically for ${\rm Re}(\rho)=-1/2$. The formal solution $\bar{\phi}_{1}(z)$ has no finite Bargmann norm for $|\alpha|<1$, in accord with the result in maciejewski2015 . There are no isolated eigenvalues in the energy range $-\Delta-2g^{2}/\omega<E<-\Delta$. Now we have two options:
@@
To decide between these two possibilities, it is not enough to state merely the non-normalizability of the entire formal solutions of (26) as in maciejewski2015 because these formal solutions are entire (and non-normalizable) also in the gaps between the discrete eigenvalues for $|\alpha|>1$.
To show that there is indeed a continuum, we must demonstrate that the formal eigenfunctions $\bar{\phi}_{\Lambda}(z)$ for eigenvalue $\Lambda$ furnish a spectral measure on the real line absolutely continuous with respect to Lebesgue measure reed-1 . For example, the formal eigenfunctions $\phi_{p}(x)=(2\pi)^{-1/2}e^{ipx}$ of the momentum operator $-i{\rm{d}}/{\rm{d}}x$ satisfy the generalized orthonormality relation,
@@
and furnish thus a proper spectral measure on the real line $-\infty<p<\infty$, because any element of $L^{2}(\mathbb{R})$ can be expanded in terms of the set $\{\phi_{p}(x)\}$ as a Fourier integral. In ${\cal B}$, the functions $\phi_{p}$ read braak2023 ,
@@
$\langle\bar{\phi}_{\Lambda_{1}}|\bar{\phi}_{\Lambda_{2}}\rangle$ $\propto\int_{S_{0}}{\rm{d}}x{\rm{d}}y\ e^{-x^{2}-y^{2}}e^{\frac{1}{2}\Big{(}(x+iy)^{2}+(x-iy)^{2}\Big{)}}(x-iy)^{\rho^{\ast}(\Lambda_{1})}(x+iy)^{\rho(\Lambda_{2})}$ (42)
@@
$=C\int_{-\infty}^{\infty}{\rm{d}}t\ \exp\left(it\left[\frac{\Lambda_{1}}{\sqrt{1-\alpha_{1}^{2}}}-\frac{\Lambda_{2}}{\sqrt{1-\alpha_{2}^{2}}}\right]\right)$ (43)
@@
with $a=\textrm{ch}\theta-\textrm{sh}\theta$ and $\textrm{ch}\theta=\alpha$. It follows with $a<1$ that $U$ is an element of $SU(1,1)$, a real subgroup of $SL_{2}(\mathbb{C})$ which acts on ${\cal B}$ as isometries duan2022 ; braak2023 . This means that there is a unitary transformation in ${\cal B}$ which maps the left hand side of (29) onto $\textrm{sh}\theta K_{0}$ which is the Hamiltonian of the harmonic oscillator with frequency $\textrm{sh}\theta$ (including the vacuum term). It follows at once that the eigenvalue equation (29) is equivalent to
@@
in accord with maciejewski2015 ; xie2019 . The energy $E$ is given implicitly in (46) and leads to a lower bound for the discrete spectrum with $\alpha>1$, namely $E>$max$(-\Delta,\Delta-\omega)$. In the case $\alpha<-1$, the same reasoning must be slightly modified because now $\textrm{ch}\theta=-\alpha$ and the transformation $U$ which yields the diagonal form in (45) has now $a=\textrm{sh}\theta+\textrm{ch}\theta>1$ and is not an element of $SU(1,1)$. If one exchanges the two eigenvectors, one obtains
@@
again in accord with maciejewski2015 ; xie2019 . Beyond the coupling regime which supports a lower discrete spectrum, the continuum threshold at $E_{thr}=-(\Delta+2g^{2}/\omega)$ is at the same time the ground state energy. As $E_{thr}<E_{c}$, the lower limit of the continuum deduced from the pole structure of the $G$-function, we see that the confluence process going from (23) to (26) is consistent with the continuum predicted from $G_{\pm}(x)$, at least regarding the energy region $E<-\Delta$. On the other hand, the $G$-function also predicts that the continuum should extend over the whole real axis from $E_{c}$ upwards. How can this be reconciled with the confluence process described in (23) and (26)?
@@
This representation preserves the canonical commutation relations. We set now $\zeta=\sqrt{2C}/\epsilon$, this means that $\zeta$ diverges as $\epsilon\rightarrow 0$. In this limit, the transformation is merely “pseudo-canonical”, a phenomenon known from systems with infinitely many degrees of freedom segal1970 . The constant $C$ has the dimension of energy. Performing the limit yields
@@
For any finite energy $E>E_{thr}$, the generalized eigenstate will be centered around $q_{\pm}(E)=\pm\zeta\sqrt{(E-E_{thr})/C}$ which diverges like $1/\epsilon=(\omega-\gamma)^{-1/2}$. The same is true of the two associated spectral measures on any interval $[E_{1},E_{2}]$ with $E_{1}>E_{thr}$ which corresponds to elements of $L^{2}(\mathbb{R})$ supported respectively on the intervals ±$\pm[q_{+}(E_{1}),q_{+}(E_{2})]$. As the normalizable states of (46) have wave functions exponentially decreasing for $q\rightarrow\pm\infty$, we conclude that they will no longer hybridize with the states in (59) for $\gamma\rightarrow\omega_{-}$ and belong to a dynamically invariant subspace of $L^{2}(\mathbb{R})$. Therefore, they can be considered as “bound states in the continuum” at the critical point. However, for all $\gamma<\omega$ the precursors of the BICs will hybridize with the precursors of the states in (59), leading to avoided crossings in the spectral graph which become more and more narrow as $\gamma$ approaches $\omega$. True BICs exists only at the critical point itself, which is here defined as the limit $\gamma\rightarrow\omega_{-}$, not by setting $\gamma=\omega$ in (9) and (10). The latter case was studied in maciejewski2015 ; xie2019 and leads just to the two discrete spectra (46), (48) and the “small continuum”, obtained by the first confluence process given in (26). The canonical transformation (54) can be implemented unitarily only if $\zeta<\infty$, so it fails exactly at $\gamma=\omega$ which is a singular point of the transformation. This phenomenon bears close resemblance to the failure to implement the scale transformation (54) unitarily with $\zeta<\infty$ but for infinitely many bosonic variables segal1970 . The Rabi-Stark model at the critical point provides thus an analogue of the infrared divergences occurring in interacting quantum field theories – just like the singular limit $\Delta/\omega\rightarrow\infty$ leads to a quantum phase transition of the QRM bakemeier2012 ; ashhab2013 ; hwang2015 , although the model contains only two degrees of freedom. The crucial difference being that in the latter case the model parameters have to take unphysical (infinite) values, whereas the Rabi-Stark model can be simulated within cavity QED right at the critical point $\gamma=\omega$ grimsmo2013 . Our results entail that this critical point has to be regularized by setting $\gamma=\omega-\epsilon^{2}$, before letting $\epsilon\rightarrow 0$ to obtain the full spectrum at $\gamma=\omega$.
IV Numerical results for large-dimensional state space